What is regression analysis? Explain steps of performing regression 

analysis in detail? 

The following data shows the yearly sales (in million Rs.) of A2Zcorporation for the last nine years. Develop a regression model. 

Year 2011, 2012, 2013, 2014, 2015, 2016, 2017, 2018 ,2019

Sale 2.3, 5.3, 5.1, 3.5, 3.4, 2.7, 2.8, 4.1, 2.9


1
Expert's answer
2022-03-01T08:28:34-0500

Solution:

Regression analysis is a powerful statistical method that allows you to examine the relationship between two or more variables of interest.

Steps in Regression Analysis

  1. Problem definition
  2. Analyse data
  3. Model Selection
  4. Model Fitting
  5. Model evaluation
  6. Conclusion


Treat 2011 as year 1, then 2019 would be 9 etc.



Sum of X = 45

Sum of Y = 32.1

Mean X = 5

Mean Y = 3.5667

Sum of squares (SSX) = 60

Sum of products (SP) = -6.6

Regression Equation = ŷ = bX + a

b = SP/SSX = -6.6/60 = -0.11

a = MY - bMX = 3.57 - (-0.11"\times" 5) = 4.11667

ŷ = -0.11X + 4.11667